OpenStudy (anonymous):

How to write in terms of sinx and/or cosx and simplify for : 1. (secx-cosx)/(secx+cosx) 2.(secx-1)/(xsecx) 3.(1-tanx)/(sinx-cosx)

hartnn (hartnn):

sec x is given as \[\sec x =\frac{1}{\cos x}\] tan x is given as \[\tan x = \frac{\sin x}{\cos x}\] that would be 1st part...

OpenStudy (anonymous):

okay

hartnn (hartnn):

so when u simplify for 1st question,are u getting \[1-\cos^2 x\] in the numerator?

hartnn (hartnn):

and \[1+\cos^2 x\] in the denominator??

OpenStudy (anonymous):

yeah, then do I make the numerator sin^2X?

hartnn (hartnn):

yes,correct,as cos^2x + sin^x =1

hartnn (hartnn):

go for 2nd one and tell what u get?

OpenStudy (anonymous):

I get (1-cosx)/(cosx^2), is that right?

hartnn (hartnn):

nopes, numerator is correct... where did x go in the denominator?? u should get: \[\frac{1-\cos x}{x}\] as final answer

OpenStudy (anonymous):

oh, right... I multiplied them wrong at the denominator

hartnn (hartnn):

and 3rd one?

OpenStudy (anonymous):

Do I make tanx into sinx/cosx for the numerator first?

hartnn (hartnn):

yes,do that

OpenStudy (anonymous):

then multiply top and bottom by cosx?

hartnn (hartnn):

yup,correct,go on....is something getting cancelled from numerator and denominator?? remember,u can write (cos x -sin x) as -(sin x -cos x)

OpenStudy (anonymous):

oh, I see. So the final will be -(1/cosx)?

hartnn (hartnn):

yes:) thats correct....its in terms of cos x if answer is allowed in sec x then it would be -sec x

OpenStudy (anonymous):

oh, thank you very much for helping!:) I appreciate it!

hartnn (hartnn):

your welcome :)